# WVSC: Wireless Video Semantic Communication with Multi-frame Compensation

### v1

@@ Abstract
Existing wireless video transmission schemes directly conduct video coding in pixel level, while neglecting the inner semantics contained in videos. In this paper, we propose a wireless video semantic communication framework, abbreviated as WVSC, which integrates the idea of semantic communication into wireless video transmission scenarios. WVSC first encodes original video frames as semantic frames and then conducts video coding based on such compact representations, enabling the video coding in semantic level rather than pixel level. Moreover, to further reduce the communication overhead, a reference semantic frame is introduced to substitute motion vectors of each frame in common video coding methods. At the receiver, multi-frame compensation (MFC) is proposed to produce compensated current semantic frame with a multi-frame fusion attention module. With both the reference frame transmission and MFC, the bandwidth efficiency improves with satisfying video transmission performance. Experimental results verify the performance gain of WVSC over other DL-based methods e.g. DVSC about 1 dB and traditional schemes about 2 dB in terms of PSNR.
Index Terms:
wireless video transmission, semantic communication, video coding, deep learning, frame compensation
@@ I Introduction
When it comes to the wireless video transmission, the key challenge is how to exploit both intra-frame correlation and inter-frame correlation to efficiently compress original videos while ensuring accurate data transmission simultaneously. Wireless video transmission schemes commonly adopt standardized video codecs such as H.264 [1] or H.265 [2] for source coding, separately followed by the channel codec such as low density parity check (LDPC). These separated source-channel coding (SSCC) schemes are obviously not optimal and would experience serious cliff effect. To address this, Deep learning (DL)-based schemes, especially semantic communications, serve as promising solutions for innovating existing SSCC-based wireless video transmission frameworks. Existing works [3, 4] have successfully conducted DL techniques into wireless video transmission designs. For example, Jiang et al. [3] have proposed a wireless semantic video conferencing system to ensure the keypoint transmission in video conference. Niu et al. [4] have proposed an end-to-end DL-enabled video semantic communication system which conducts SNR-adaptive channel codec along with semantic repairment to combat multi-dimensional noise.
@@
Notations: $\mathbb{R}$ and $\mathbb{C}$ refer to the real and complex number sets, respectively. $\mathcal{CN}\left(\mu,\sigma^{2}\right)$ denotes a complex Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. $\mathrm{diag}(\cdot)$ refer to the diagonalization operations between a vector and its corresponding diagonal matrix. ||$\cdot$|| ⋅ | refers to computing the modulus of a complex number. $\ast$ refers to the element-wise multiplication. $\mathbf{I}$ denotes the unit matrix. The operator $\left(\cdot\right)^{T}$ denotes the matrix transpose.
@@ II-A Semantic I frame transmission
We consider a GoP-based wireless video transmission system. Given a video sequence $\mathcal{X}=\left\{\mathbf{X}_{1},\mathbf{X}_{2},\cdots,\mathbf{X}_{N}\right\}$, where $\mathbf{X}_{n}\in\mathbb{R}^{I\times H\times W\times 3},n=1,\cdots,N$, which is composed of $N$ successive GoPs. For an arbitrary GoP $\mathbf{X}=\left\{\mathbf{x}^{1},\mathbf{x}^{2},\cdots,\mathbf{x}^{I}\right\}\in\mathcal{X}$, where $I$ is the total number of frames in a GoP, $\mathbf{x}^{i}\in\mathbb{R}^{H\times W\times 3}$, we treat the first frame $\mathbf{x}^{1}$ as the reference frame $\mathbf{x}^{\mathrm{ref}}$. The reference frame, which is also the I frame in a GoP, is communicated using DL-based JSCC, i.e., the semantic encoder and decoder in Fig. 2, similar to the simple image semantic communications. The semantic encoder, $f_{e}(\cdot):\mathbb{R}^{I\times H\times W\times 3}\longrightarrow\mathbb{R}^{I\times L}$, encodes $\mathcal{X}$ into the semantic sequence, =$\mathcal{F}=\left\{\mathbf{f}^{1},\mathbf{f}^{2},\cdots,\mathbf{f}^{I}\right\}$, where $L$ is the code length for the extracted semantic frame with $\mathbf{f}^{i}\in\mathbb{R}^{L}$. Here $\mathbf{f}^{\mathrm{ref}}=\mathbf{f}^{1}$ is the semantic I frame. Then, the semantic I frame is transmitted through wireless channels and equalized by the minimum mean square error (MMSE) channel equalization, which can be formulated as
@@
At the receiver, the semantic decoder, $g_{a}(\cdot):\mathbb{R}^{I\times L}\longrightarrow\mathbb{R}^{I\times H\times W\times 3}$, converts ^=$\hat{\mathcal{F}}=\left\{\hat{\mathbf{f}^{1}},\hat{\mathbf{f}^{2}},\cdots,\hat{\mathbf{f}^{I}}\right\}$ into the final reconstructed video sequence $\hat{\mathcal{X}}=\left\{\hat{\mathbf{X}}^{1},\hat{\mathbf{X}}^{2},\cdots,\hat{\mathbf{X}}^{I}\right\}$. The whole transmission process for the semantic I frame can be formulated as
@@ II-B Semantic P frame transmission
For the semantic P frame transmission, non-reference semantics $\mathbf{f}^{i}$ with $i\geq 2$, are communicated using not only DL-based JSCC but also additional pre-/post-processing. They are processed by the motion estimation &\&& compensation video encoder, $V_{c_{1}}(\cdot,\mathbf{f}^{\mathrm{ref}}):\mathbb{R}^{L}\longrightarrow\mathbb{R}^{L}$, along with the reference semantics $\mathbf{f}^{\mathrm{ref}}$ to acquire the predicted semantic frame $\bar{\mathbf{f}}^{i}\in\mathbb{R}^{L}$. Then, the residual $\mathbf{r}^{i}\in\mathbb{R}^{L}$ can be computed through the subtraction between $\mathbf{f}^{i}$ and $\bar{\mathbf{f}}^{i}\in\mathbb{R}^{L}$. In this way, the transmission for the P frame semantics are substituted by the corresponding sparse residuals which is much easier to be compressed than original video frames. Notably, in common video coding schemes, $\bar{\mathbf{f}}^{i}$ is known for both compression and decompression process. However, for wireless video transmission scenarios, $\bar{\mathbf{f}}^{i}$ is obviously unknown at the receiving end. Instead of transmitting motion vectors for every semantic P frame to recover $\bar{\mathbf{f}}^{i}$, we utilize the consistent reference semantic frame $\mathbf{f}^{\mathrm{ref}}$ as the substitute for the motion vectors, greatly saving the communication cost. After that, the residuals are further compressed by the residual encoder, $h_{s}(\cdot):\mathbb{R}^{L}\longrightarrow\mathbb{R}^{L_{1}}$, to provide adjustable transmission rates for the semantic I frame and compressed residuals $\mathbf{r}_{\mathrm{c}}^{i}$, commonly preallocating more transmission rate for the semantic I frame to ensure the latter deep video decoding performances at the receiver. In this way, the semantic P frame is converted to the compactly compressed residuals for transmission. For simplicity, the transmitted $\mathbf{r}_{\mathrm{c}}^{i}$ in a GoP share the same channel states with the semantic I frame. The wireless transmission process for the compressed residuals of semantic P frames can be formulated as
@@
At the receiver, the residual decoder, $h_{p}(\cdot):\mathbb{R}^{L_{1}}\longrightarrow\mathbb{R}^{L}$, recovers the residuals as $\hat{\mathbf{r}}^{i}\in\mathbb{R}^{L}$. With the received $\hat{\mathbf{f}}^{\mathrm{ref}}$ and $\hat{\mathbf{r}}^{i}$, we obtain the reconstructed semantic sequence, ^=$\hat{\mathcal{F}}=\left\{\hat{\mathbf{f}}^{1},\hat{\mathbf{f}}^{2},\cdots,\hat{\mathbf{f}}^{I}\right\}$, using a series of frame compensation and video decoding operations. The multi-frame compensation module, $C(\cdot,\mathbf{f},\mathbf{h}):\mathbb{R}^{L}\longrightarrow\mathbb{R}^{L}$, is employed to polish $\mathbf{\hat{f}}^{\mathrm{ref}}$ into the current semantic P frame $\mathbf{f}^{i}$ as much as possible, where $\mathbf{f}=\left\{\mathbf{\hat{f}}^{i-1},\cdots,\mathbf{\hat{f}}^{2}\right\}$ refers to the previously reconstructed frames. $\mathbf{\tilde{f}}^{i}\in\mathbb{R}^{L}$ refers to the polished current semantic frame after multi-frame compensation. After multi-frame compensation, through motion estimation &\&& compensation video decoder $V_{c_{2}}(\cdot,\mathbf{\hat{f}}^{\mathrm{ref}}):\mathbb{R}^{L}\longrightarrow\mathbb{R}^{L}$, the reconstructed predicted semantic frame $\mathbf{\check{f}}^{i}\in\mathbb{R}^{L}$ can be acquired. With $\hat{\mathbf{r}}^{i}$ and $\mathbf{\check{f}}^{i}$, we are able to acquire the current reconstructed semantic frames $\mathbf{\hat{f}}^{i}$ by adding them together. Finally, the semantic decoder decodes $\mathbf{\hat{f}}^{i}$ into final reconstructed video frames $\mathbf{\hat{x}}^{i}$. The whole process can be formulated as
@@ III-A Deep Video Coding in Semantic Level
Figure 3: The structure of motion estimation &\&& compensation network.
As illustrated in Section II, WVSC conducts deep video coding in semantic level. That is to say, the inter-frame correlation is exploited based on the semantics extracted by original frames. The structure of motion estimation &\&& compensation video coder is shown in Fig. 3. It is mentioned that the video encoder ${V_{c_{1}}(\cdot,\mathbf{f}^{\mathrm{ref}})}$ and decoder $V_{c_{2}}(\cdot,\mathbf{\hat{f}}^{\mathrm{ref}})$ share the same network structure. The video coder is composed of two parts: motion estimation and motion compensation. For the motion estimation, the current semantic frame $\mathbf{f}^{i}$ and the reference semantic frame $\mathbf{f}^{\mathrm{ref}}$ are concatenated together as input. Then through a lightweight structure composed of a series of convolutional layers and activation layers with LeakyReLU [6], the motion offset map $O_{i}$ between reference and current semantics is generated. In other words, $O_{i}$ refers to the motion vector in common video coding frameworks. For the motion compensation, with the estimated offset map, the predicted semantic frame $\mathbf{\bar{f}}^{i}$ can be produced by the motion compensation network which is based on the residual convolutional networks.
@@ IV-A1 Datasets
For the wireless video semantic transmission, we quantify the performances of proposed WVSC versus other benchmarks over the UCF101 [8] dataset, which is a widely-used dataset for deep video compression. The dataset is splited into about 5:1 ratio for training and testing, respectively. During model training, images are randomly cropped to 128$\times$128$\times$3.
@@ IV-A2 Model Deployment Details
The network deployment of WVSC utilizes the Swin-Transformer [9] backbone as the semantic codec with $\{N_{1},N_{2},N_{3},N_{4}\}=\{2,2,2,2\}$ Transformer blocks. Various signal-to-noise ratios (SNRs) are randomly sampled for each image batch during training stage to acquire SNR-adaptive results in various test conditions. For model training, we use variable learning rate, which decreases step-by-step from 1e-4 to 2e-5. The batchsize is set as 1. The default GoP size $N$ is 5. The whole framework is optimized with Adam [10] algorithm. All the experiments of WVSC and other DL-based benchmarks are runned in RTX3090 GPUs with Pytorch.
@@ IV-A3 Comparison Benchmarks
H.264+LDPC+QAM: The separated source and channel coding scheme with H.264 [1] video codec as source coding and Low-density Parity-check (LDPC) as channel coding, along with the quadrature amplitude modulation (QAM).
H.265+LDPC+QAM: The separated source and channel coding scheme with H.265 [2] video codec as source coding and LDPC as channel coding, along with the QAM.
(a) PNSR for the reconstructed videos.
(b) MS-SSIM for the reconstructed videos.
(c) LPIPS for the reconstructed videos.
Figure 5: Quality of the reconstructed images versus the SNRs under Rayleigh fading channels (CBR = 0.10).

### v2

@@ Abstract
Existing wireless video transmission schemes directly conduct video coding in pixel level, while neglecting the inner semantics contained in videos. In this paper, we propose a wireless video semantic communication framework, abbreviated as WVSC, which integrates the idea of semantic communication into wireless video transmission scenarios. WVSC first encodes original video frames as semantic frames and then conducts video coding based on such compact representations, enabling the video coding in semantic level rather than pixel level. Moreover, to further reduce the communication overhead, a reference semantic frame is introduced to substitute motion vectors of each frame in common video coding methods. At the receiver, multi-frame compensation (MFC) is proposed to produce compensated current semantic frame with a multi-frame fusion attention module. With both the reference frame transmission and MFC, the bandwidth efficiency improves with satisfying video transmission performance. Experimental results verify the performance gain of WVSC over other DL-based methods (e.g. DVSC) about 1 dB and traditional standardized schemes (e.g. H.265+5G LDPC) about 0.7 dB in terms of PSNR.
@@ I Introduction
When it comes to the wireless video transmission, the key challenge is how to exploit both intra-frame correlation and inter-frame correlation to efficiently compress original videos while ensuring accurate data transmission simultaneously. Wireless video transmission schemes commonly adopt standardized video codecs such as H.264 [1] or H.265 [2] for source coding, separately followed by the channel codec such as low density parity check (LDPC). These separated source-channel coding (SSCC) schemes are obviously not optimal and would experience serious cliff effect. To address this, Deep learning (DL)-based schemes, especially semantic communications [3, 4, 5], serve as promising solutions for innovating existing SSCC-based wireless video transmission frameworks. Existing works [4, 5] have successfully conducted DL techniques into wireless video transmission designs. For example, Jiang et al. [4] have proposed a wireless semantic video conferencing system to ensure the keypoint transmission in video conference. Niu et al. [5] have proposed an end-to-end DL-enabled video semantic communication system which conducts SNR-adaptive channel codec along with semantic repairment to combat multi-dimensional noise.
@@
Notations: $\mathbb{R}$ and $\mathbb{C}$ refer to the real and complex number sets, respectively. $\mathcal{CN}\left(\mu,\sigma^{2}\right)$ denotes a complex Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. $\mathrm{diag}(\cdot)$ refer to the diagonalization operations between a vector and its corresponding diagonal matrix. $|\cdot|$ refers to computing the modulus of a complex number. $\ast$ refers to the element-wise multiplication. $\mathbf{I}$ denotes the unit matrix. The operator $\left(\cdot\right)^{T}$ denotes the matrix transpose.
@@ II-A Semantic I frame transmission
We consider a GoP-based wireless video transmission system. Given a video sequence $\mathcal{X}=\left\{\mathbf{X}_{1},\mathbf{X}_{2},\cdots,\mathbf{X}_{N}\right\}$, where $\mathbf{X}_{n}\in\mathbb{R}^{I\times H\times W\times 3},n=1,\cdots,N$, which is composed of $N$ successive GoPs. For an arbitrary GoP $\mathbf{X}=\left\{\mathbf{x}^{1},\mathbf{x}^{2},\cdots,\mathbf{x}^{I}\right\}\in\mathcal{X}$, where $I$ is the total number of frames in a GoP, $\mathbf{x}^{i}\in\mathbb{R}^{H\times W\times 3}$, we treat the first frame $\mathbf{x}^{1}$ as the reference frame $\mathbf{x}^{\mathrm{ref}}$. The reference frame, which is also the I frame in a GoP, is communicated using DL-based JSCC, i.e., the semantic encoder and decoder in Fig. 2, similar to the simple image semantic communications. The semantic encoder, $f_{e}(\cdot):\mathbb{R}^{I\times H\times W\times 3}\longrightarrow\mathbb{R}^{I\times L}$, encodes $\mathcal{X}$ into the semantic sequence, $\mathcal{F}=\left\{\mathbf{f}^{1},\mathbf{f}^{2},\cdots,\mathbf{f}^{I}\right\}$, where $L$ is the code length for the extracted semantic frame with $\mathbf{f}^{i}\in\mathbb{R}^{L}$. Here $\mathbf{f}^{\mathrm{ref}}=\mathbf{f}^{1}$ is the semantic I frame. Then, the semantic I frame is transmitted through wireless channels and equalized by the minimum mean square error (MMSE) channel equalization. It can be formulated as
@@
At the receiver, the semantic decoder, $g_{a}(\cdot):\mathbb{R}^{I\times L}\longrightarrow\mathbb{R}^{I\times H\times W\times 3}$, converts $\hat{\mathcal{F}}=\left\{\hat{\mathbf{f}^{1}},\hat{\mathbf{f}^{2}},\cdots,\hat{\mathbf{f}^{I}}\right\}$ into the final reconstructed video sequence $\hat{\mathcal{X}}=\left\{\hat{\mathbf{X}}^{1},\hat{\mathbf{X}}^{2},\cdots,\hat{\mathbf{X}}^{I}\right\}$. The whole transmission process for the semantic I frame can be formulated as
@@ II-B Semantic P frame transmission
For the semantic P frame transmission, non-reference semantics $\mathbf{f}^{i}$ with $i\geq 2$, are communicated using not only DL-based JSCC but also additional pre-/post-processing. They are processed by the motion estimation &\& compensation video encoder, $V_{c_{1}}(\cdot,\mathbf{f}^{\mathrm{ref}}):\mathbb{R}^{L}\longrightarrow\mathbb{R}^{L}$, along with the reference semantics $\mathbf{f}^{\mathrm{ref}}$ to acquire the predicted semantic frame $\bar{\mathbf{f}}^{i}\in\mathbb{R}^{L}$. Then, the residual $\mathbf{r}^{i}\in\mathbb{R}^{L}$ can be computed through the subtraction between $\mathbf{f}^{i}$ and $\bar{\mathbf{f}}^{i}\in\mathbb{R}^{L}$. In this way, the transmission for the P frame semantics are substituted by the corresponding sparse residuals which is much easier to be compressed than original video frames. Notably, in common video coding schemes, $\bar{\mathbf{f}}^{i}$ is known for both compression and decompression process. However, for wireless video transmission scenarios, $\bar{\mathbf{f}}^{i}$ is obviously unknown at the receiving end. Instead of transmitting motion vectors for every semantic P frame to recover $\bar{\mathbf{f}}^{i}$, we utilize the consistent reference semantic frame $\mathbf{f}^{\mathrm{ref}}$ as the substitute for the motion vectors, greatly saving the communication cost. After that, the residuals are further compressed by the residual encoder, $h_{s}(\cdot):\mathbb{R}^{L}\longrightarrow\mathbb{R}^{L_{1}}$, to provide adjustable transmission rates for the semantic I frame and compressed residuals $\mathbf{r}_{\mathrm{c}}^{i}$, commonly preallocating more transmission rate for the semantic I frame to ensure the latter deep video decoding performances at the receiver. In this way, the semantic P frame is converted to the compactly compressed residuals for transmission. For simplicity, the transmitted $\mathbf{r}_{\mathrm{c}}^{i}$ in a GoP share the same channel states with the semantic I frame. The wireless transmission process for the compressed residuals of semantic P frames can be formulated as
@@
At the receiver, the residual decoder, $h_{p}(\cdot):\mathbb{R}^{L_{1}}\longrightarrow\mathbb{R}^{L}$, recovers the residuals as $\hat{\mathbf{r}}^{i}\in\mathbb{R}^{L}$. With the received $\hat{\mathbf{f}}^{\mathrm{ref}}$ and $\hat{\mathbf{r}}^{i}$, we obtain the reconstructed semantic sequence, $\hat{\mathcal{F}}=\left\{\hat{\mathbf{f}}^{1},\hat{\mathbf{f}}^{2},\cdots,\hat{\mathbf{f}}^{I}\right\}$, using a series of frame compensation and video decoding operations. The multi-frame compensation module, $C(\cdot,\mathbf{f},\mathbf{h}):\mathbb{R}^{L}\longrightarrow\mathbb{R}^{L}$, is employed to polish $\mathbf{\hat{f}}^{\mathrm{ref}}$ into the current semantic P frame $\mathbf{f}^{i}$ as much as possible, where $\mathbf{f}=\left\{\mathbf{\hat{f}}^{i-1},\cdots,\mathbf{\hat{f}}^{2}\right\}$ refers to the previously reconstructed frames. $\mathbf{\tilde{f}}^{i}\in\mathbb{R}^{L}$ refers to the polished current semantic frame after multi-frame compensation. After multi-frame compensation, through motion estimation &\& compensation video decoder $V_{c_{2}}(\cdot,\mathbf{\hat{f}}^{\mathrm{ref}}):\mathbb{R}^{L}\longrightarrow\mathbb{R}^{L}$, the reconstructed predicted semantic frame $\mathbf{\check{f}}^{i}\in\mathbb{R}^{L}$ can be acquired. With $\hat{\mathbf{r}}^{i}$ and $\mathbf{\check{f}}^{i}$, we are able to acquire the current reconstructed semantic frames $\mathbf{\hat{f}}^{i}$ by adding them together. Finally, the semantic decoder decodes $\mathbf{\hat{f}}^{i}$ into final reconstructed video frames $\mathbf{\hat{x}}^{i}$. The whole process can be formulated as
@@ III-A Deep Video Coding in Semantic Level
Figure 3: The structure of motion estimation &\& compensation network.
As illustrated in Section II, WVSC conducts deep video coding in semantic level. That is to say, the inter-frame correlation is exploited based on the semantics extracted by original frames. The structure of motion estimation &\& compensation video coder is shown in Fig. 3. It is mentioned that the video encoder $V_{c_{1}}(\cdot,\mathbf{f}^{\mathrm{ref}})$} and decoder $V_{c_{2}}(\cdot,\mathbf{\hat{f}}^{\mathrm{ref}})$ share the same network structure. The video coder is composed of two parts: motion estimation and motion compensation. For the motion estimation, the current semantic frame $\mathbf{f}^{i}$ and the reference semantic frame $\mathbf{f}^{\mathrm{ref}}$ are concatenated together as input. Then through a lightweight structure composed of a series of convolutional layers and activation layers with LeakyReLU [7], the motion offset map $O_{i}$ between reference and current semantics is generated. In other words, $O_{i}$ refers to the motion vector in common video coding frameworks. For the motion compensation, with the estimated offset map, the predicted semantic frame $\mathbf{\bar{f}}^{i}$ can be produced by the motion compensation network which is based on the residual convolutional networks.
@@ III-B Multi-frame Compensation for the Reference Frame
III-C Correlation between the I frame and the P frame
For the WVSC, I frames and P frames have distinct transmission characteristics. Since the I frame serves as the key reference for the subsequent P frames in a GoP, more bandwidth resources should be allocated to the semantic I frame than to the semantic P frames. In this way, the residual encoder and decoder act as a controller to adjust the allocated bitrate between the semantic I frame and subsequent semantic P frames, thereby providing unequal error protection for the overall wireless video transmission.
The WVSC is trained using a successive training strategy, meaning it is optimized not only within a single frame but also across different GoPs to jointly train the I frames and P frames. The end-to-end video reconstruction loss is formulated as follows
[…]
where mean square error (MSE) is set as the default loss, $\hat{\mathbf{x}}_{n}^{i}$ and ${\mathbf{x}}_{n}^{i}$ refer to the $i$-th reconstructed video frame and original frame in the $n$-th GoP, respectively.
@@ IV-A1 Datasets
For the wireless video semantic transmission, we quantify the performances of proposed WVSC versus other benchmarks over the UCF101 [9] dataset, which is a widely-used dataset for deep video compression. The dataset is splited into about 5:1 ratio for training and testing, respectively. During model training, images are randomly cropped to $\times128$$\times$3.
(a) PSNR for the reconstructed videos.
(b) MS-SSIM for the reconstructed videos.
(c) LPIPS for the reconstructed videos.
Figure 5: Quality of the reconstructed images versus the SNRs under Rayleigh fading channels (CBR = 0.04).
(a) PSNR for the reconstructed videos.
(b) MS-SSIM for the reconstructed videos.
(c) LPIPS for the reconstructed videos.
@@ IV-A2 Model Deployment Details
The network deployment of WVSC utilizes the Swin-Transformer [10] backbone as the semantic codec with $\{N_{1},N_{2},N_{3},N_{4}\}=\{2,2,2,2\}$ Transformer blocks. Various signal-to-noise ratios (SNRs) are randomly sampled for each image batch during training stage to acquire SNR-adaptive results in various test conditions. For model training, we use variable learning rate, which decreases step-by-step from 1e-4 to 2e-5. The batchsize is set as 1. The default GoP size $N$ is 10. The whole framework is optimized with Adam [11] algorithm. All the experiments of WVSC and other DL-based benchmarks are conducted in RTX3090 GPUs with Pytorch2.0.0.
@@ IV-A3 Comparison Benchmarks
H.264+LDPC+QAM: The separated source and channel coding scheme with H.264 [1] video codec as source coding with GoP size 32 and 5G Low-density Parity-check (LDPC) [12] as channel coding, along with the quadrature amplitude modulation (QAM).
H.265+LDPC+QAM: The separated source and channel coding scheme with H.265 [2] video codec as source coding with GoP size 32 and 5G LDPC as channel coding, along with the QAM.
Note that DVSC is the DL-based pixel-level point-to-point wireless transmission scheme, which corresponds to Fig. 1(a). H.265/H.264+LDPC+QAM are SSCC schemes which H.265/H.264 are employed by ffmpeg-python with low delay pattern while 5G LDPC is implemented by [12] with code length 4096.